# Thompson sampling for zero-inflated count outcomes with an application to the Drink Less mobile health study

### v1

@@ 5 Theoretical Results
$\mathcal{R}(T,\gamma^{*},\beta^{*})$$\leq$ $q(c_{1}^{\beta}+c_{1}^{\gamma}+c_{2}^{\beta}+c_{2}^{\gamma})\left(1+\frac{2}{0.0225-2/T+1/T^{2}}\right)\times\sqrt{2dT\log(T/d)}$
$+$ $(\tau+4-2/T)\Delta_{\max},$
where
$c_{1}^{\beta}$ $=$ $Y_{\max}\kappa_{1,\min}^{-1}\sqrt{d\log(T/d)+2\log T},$
$c_{1}^{\gamma}$ $=$ $\kappa_{2,\min}^{-1}\sqrt{d\log(T/d)+2\log T},$
$c_{2}^{\beta}$ $=$ $c_{1}^{\beta}\sqrt{2\kappa_{1,\min}^{-1}\kappa_{1,\max}\log(KT)},$
$c_{2}^{\gamma}$ $=$ $c_{1}^{\gamma}\sqrt{2\kappa_{2,\min}^{-1}\kappa_{2,\max}\log(KT)}.$
Theorem 4 also establishes $\tilde{O}(d\sqrt{T\log K})$ regret for TS-ZIP. A comprehensive proof can be found in Section B.2 of the Supplement. Notably, the frequentist regret for TS-ZINB aligns with Equation (4), albeit under slightly different conditions, as detailed in Section B.2 of the Supplement.
TS-Count with normal approximation starts with an initial exploration time period $\tau$, which consists of delivering interventions $A_{1},A_{2},\cdots,A_{\tau}$ with uniform randomization probabilities. This can ensure the invertibility of $G_{t}$ for all $t\geq\tau$. The necessity is motivated by the classical likelihood theory: as the sample size $t\rightarrow\infty$, the MLE $\hat{\beta}_{t}$ is consistent and asymptotically normal, i.e., $\hat{\beta}_{t}-\beta^{*}\rightarrow\mathcal{N}(0,\mathcal{I}_{t}^{-1}(\beta^{*}))$. The invertibility of $G_{t}$ can ensure that $\mathcal{I}_{t}$ is invertible and a unique MLE is attainable. Proposition 1 provides a way for theoretically specifying the length of the initial exploration time.
Proposition 1.
This is an adaptation of Proposition 1 in Li et al. (2017) and Oh and Iyengar (2019) to TS-Count, and the proof can be found in the same sources. Note that the i.i.d. assumption on $X_{t}$ is only required during the initial exploration period. Subsequently, $X_{t}$ can be chosen adversarially, without assuming any structure on it, as long as $||X_{t}||$ is bounded. In terms of the Drink Less app, Proposition 1 provides guidance on the number of historical observations required to ensure the validity of the learning algorithm for intervention delivery.
@@ 6 Simulation Studies
6.1 Data Generation
Our simulation studies encompass a total of eight settings, with the stochastic proximal outcome generated from (1) a Poisson distribution, (2) an overdispersed Poisson distribution with a low dispersion level, (3) an overdispersed Poisson distribution with a moderate dispersion level, (4) an overdispersed Poisson distribution with a high dispersion level, (5) a ZIP distribution, (6) a zero-inflated and overdispersed Poisson distribution with a low dispersion level, (7) a zero-inflated and overdispersed Poisson distribution with a moderate dispersion level, and (8) a zero-inflated and overdispersed Poisson distribution with a high dispersion level.
@@ 6.1 Data Generation
$Y_{t}\sim\operatorname{Poisson}(\lambda_{t}\mu_{t}),~{}\ \lambda_{t}\sim\operatorname{Gamma}(\omega,1/\omega),~{}\ \mu_{t}=\exp\left(\phi(A_{t},X_{t})^{\top}\beta^{*}\right).$
In this case, the outcomes are generated from an overdispersed Poisson model, where $\lambda_{t}$ introduces extra overdispersion, and $\omega$ controls the level of dispersion. We induce high, moderate, and low levels of overdispersion by setting $\omega=0.25,1,$ and $25$, respectively. When setting $\lambda_{t}=1$, we recover the Poisson distribution.
@@
$Y_{t}$ $=$ $Z_{t}C_{t}.$
Here, the outcomes are generated from a zero-inflated and overdispersed Poisson model. We also consider scenarios without zero-inflation and overdispersion by setting both $Z_{t}$ and $\lambda_{t}$ to $1$.
For the simulation setup, we consider a scenario with $K=20$ intervention options. For each intervention $a$ at every time point $t$, the feature vector $\phi(a,X_{t})$ is directly sampled from a multivariate normal distribution, specifically $\operatorname{MVN}(\mathbf{0},\mathbf{I}_{4\times 4})$, where $\mathbf{I}_{4\times 4}$ denotes the diagonal unit matrix. Moreover, we generate the true parameters $\gamma^{*}$ and $\beta^{*}$ from a multivariate normal distribution $\operatorname{MVN}(\mathbf{0},\mathbf{I}_{4\times 4})$. We normalize the feature and true parameters to satisfy $\|\phi(a,X_{t})\|\leq 1$, $\|\gamma^{*}\|\leq 1$ and $\|\beta^{*}\|\leq 1$.
For each setting, we set the time horizon to $T=1000$ and the initial random exploration time to $\tau=20$. To form the initial estimators, we selected each intervention once in the first $20$ decisions. We set the tuning parameter $\alpha=1$ for all simulations, which enables the recovery of the Laplace approximation. We conduct $200$ replications and average the results. The decision-making performance is evaluated using the observed average regret, i.e., the cumulative difference between the expected outcome of interventions selected by the algorithm and that of the optimal intervention.
@@ 6.2 Results
The results under Settings (1) - (4) are presented in Figure 2. Our findings reveal that the TS-Count variants exhibit superior performance compared to the Linear TS algorithm with even log transformation, which demonstrates consistently high and linear regret across all settings. Furthermore, our results suggest that TS-Poisson outperforms other algorithms, including TS-NB, particularly in settings with varying degrees of overdispersion. This is attributed to the reliability of Poisson regression in yielding consistent estimates when the conditional mean model is appropriately specified, meeting the requirement for sublinear regret, as demonstrated in the proof of regret.
@@ 7 Application to the Drink Less Case Study
In this section, we implement the proposed algorithms in a real case study - the Drink Less MRT - to evaluate and benchmark their performances. We consider not only the TS-Count variants and the Linear TS algorithm that incorporates a logarithmic transformation of the outcome, but also a uniform random strategy that assigns interventions at a constant and balanced probability $1/K$. The application to the Drink Less MRT posits the unique challenge of solving $N$ contextual bandit problems while ensuring valid post-study analysis, as in a typical randomized experiment. We outline our strategies for tackling these challenges, followed by a detailed description of the simulation environment and the results.
7.1 Solving Multiple Contextual Bandit Problems
@@
Up to this point, our discussion has centered on an individual user. However, within the framework of mHealth studies or MRTs, the paradigm typically involves $N$ users, each receiving interventions at $T$ decision times. To be specific, at any decision time $t$, we have a collection of historical observations denoted by $D_{n,t-1}=\left\{X_{n,i},A_{n,i},Y_{n,i};i=1,\cdots,t-1\right\}_{n=1}^{N}\cup\{X_{n,t}\}_{n=1}^{N}$. The overall objective is to customize interventions to cater to the specific demands of each user. This scenario can be envisioned as $N$ separate contextual bandit problems, where each participant is associated with a unique model parameter $\eta_{n}^{*}$. In applying our methods to the Drink Less data, we deploy the proposed algorithms on a per-user basis. This approach aligns with previous research, which has successfully employed user-centric algorithms with notable effectiveness (Liao et al., 2020; Yao et al., 2021).
7.2 Ensuring Valid Post-Study Analysis
@@
When applying the proposed algorithms to MRTs, a critical aspect is the facilitation of robust post-study statistical analyses. A central component of this analysis is the estimation of treatment effects, which provide valuable insights for future mHealth interventions. Yet, it is worth noting that an algorithm designed solely to minimize regret may not ensure error controls or sufficient statistical power for drawing reliable inferences regarding treatment effects with standard inference methods, nor may it support off-policy evaluations effectively (Liao et al., 2020; Yao et al., 2021). Specifically, the randomization probabilities must remain strictly between 00 and $1$, avoiding the extremes. In light of this, we have adapted the TS-Count algorithm by incorporating the following modifications:
  1. 1.
    Calculate the randomization probability.
    • (a)
      For TS-Poisson and TS-NB:
      $p_{a,n,t}:=\operatorname{Pr}\big{[}\phi(a,X_{n,t})^{\top}\tilde{\beta}_{n,t}=\max_{b}\phi(b,X_{n,t})^{\top}\tilde{\beta}_{n,t})\big{]},$
      for any $a\in\mathcal{A}$, where $\tilde{\beta}_{n,t}\sim\mathcal{N}(\hat{\beta}_{n,t},\alpha_{\beta}^{2}\mathcal{I}_{n,t}^{-1}(\hat{\beta}_{n,t}))$.
    • (b)
      TS-ZIP and TS-ZINB:
      $p_{a,n,t}$ $:=\operatorname{Pr}\big{[}\{1-\operatorname{sigmoid}(\phi(a,X_{n,t})^{\top}\tilde{\gamma}_{n,t})\}\exp\{\phi(a,X_{n,t})^{\top}\tilde{\beta}_{n,t}\}=$
      $\max_{b}\{1-\operatorname{sigmoid}(\phi(b,X_{n,t})^{\top}\tilde{\gamma}_{n,t})\}\exp\{\phi(b,X_{n,t})^{\top}\tilde{\beta}_{n,t}\}\big{]},$
      for any $a\in\mathcal{A}$, where $\tilde{\gamma}_{n,t}\sim\mathcal{N}(\hat{\gamma}_{n,t},\alpha_{\gamma}^{2}\mathcal{I}_{n,t}^{-1}(\hat{\gamma}_{n,t}))$ and $\tilde{\beta}_{n,t}\sim\mathcal{N}(\hat{\beta}_{n,t},\alpha_{\beta}^{2}\mathcal{I}_{n,t}^{-1}(\hat{\beta}_{n,t}))$
  2. 2.
    Constrain the randomization probability, i.e., $\tilde{p}_{a,n,t}=\min(p_{\max},\max(p_{a,n,t},p_{\min}))$, where $10<p_{\min}<p_{\max}<$ 1.
  3. 3.
    Select intervention $A_{n,t}=a$ with probability $\tilde{p}_{a,n,t}$.
In the preceding sections, we avoid explicitly computing the randomization probability for each intervention following a posterior sampling as shown in Algorithms 3-4. However, the determination and documentation of these probabilities become indispensable in the setting of MRTs for post-study analysis. These randomization probabilities are often instrumental for the accurate estimation of causal effects (Boruvka et al., 2018; Qian et al., 2021; Liu et al., 2023b). Furthermore, these probabilities are key in many novel inferential approaches specifically designed for adaptive algorithms like TS (Hadad et al., 2021; Deliu et al., 2021).
7.3 Simulation Environment
To evaluate the performances of the proposed algorithms, we redesign and analyze the Drink Less MRT, according to a simulation environment guided by real data. We show that the proposed algorithms substantially enhance the depth of user engagement with the Drink Less app, measured by the number of screen views following each intervention decision time. We detail the simulation framework as follows.
For each user $n$ at decision time $t$, after the delivery of an intervention $A_{n,t}$, the environment yields an outcome in response to the intervention using a zero-inflated overdispersed Poisson model:
$Z_{n,t}$ $\sim$ $\operatorname{Bernoulli}(1-p_{n,t}),~{}\ p_{n,t}=\operatorname{sigmoid}\left(\phi(A_{n,t},X_{n,t})^{\top}\gamma_{n}^{*}\right),$
$C_{n,t}$ $\sim$ $\operatorname{Poisson}(\lambda_{n,t}\mu_{n,t}),~{}\ \lambda_{n,t}\sim\operatorname{Gamma}(\omega,1/\omega),~{}\ \mu_{n,t}=\exp\left(\phi(A_{n,t},X_{n,t})^{\top}\beta_{n}^{*}\right),$
$Y_{n,t}$ $=$ $Z_{n,t}C_{n,t}.$
Within this model, the term $\phi(A_{n,t},X_{n,t})$ represents the feature vector, containing the administered interventions, the covariate reflecting the number of days since the app was downloaded, and the interaction effect between the interventions and the covariate. We determine the true model parameters, $\gamma_{n}^{*}$ and $\beta_{n}^{*}$, by fitting a ZINB mixed-effects model to the data from the Drink Less MRT.

### v2

@@ 5 Regret Analysis
$\mathcal{R}(T,\gamma^{*},\beta^{*})$$\leq$ $(\tau+4-2/T)\Delta_{\max}+qD\sqrt{d\log(T/d)+2\log T}\sqrt{2dT\log(T/d)}$
$\times$ $\Bigg{[}Y_{\max}\kappa_{1,\min}^{-1}\left(1+\sqrt{2\kappa_{1,\min}^{-1}\kappa_{1,\max}\log(KT)}\right)$
$+$ $\kappa_{2,\min}^{-1}\left(1+\sqrt{2\kappa_{2,\min}^{-1}\kappa_{2,\max}\log(KT)}\right)\Bigg{]},$
where the number of exploration rounds $\tau$ satisfies $\lambda_{\min}(G_{\tau})\geq 1$. The constant $D$ is in the proof.
Theorem 5.4 also establishes $\tilde{O}(d\sqrt{T\log K})$ frequentist regret bound for TS-ZIP. A comprehensive proof can be found in Section S2.2 of the Supplement. Notably, the frequentist regret bound for TS-ZINB aligns with Theorem 5.4, albeit under slightly different conditions, as detailed in Section S2.2 of the Supplement.
Remark 5.1.
The regret bound $\tilde{O}(d\sqrt{T\log K})$ aligns with the regret bounds observed in generalized linear bandits with finite interventions (Kveton et al., 2020; Filippi et al., 2010; Li, Lu and Zhou, 2017), up to a factor of $\sqrt{\log K}$. Importantly, Li, Lu and Zhou (2017) suggests that a better frequentist regret bound for TS-Count can be established by imposing an additional condition on the minimum eigenvalue of $G_{t}$, that is, $\sum_{\tau=1}^{T}\lambda_{\min}^{-1/2}(G_{t})\leq z\sqrt{T}$. The arguments presented by Li, Lu and Zhou (2017) are also applicable in this context.
@@ 6 Simulation Studies
6.1 Setup
For the simulation setup, we consider a scenario with $K=20$ intervention options over a time horizon $T=1000$. For each intervention $a$ at every time point $t$, the feature vector $\phi(a,X_{t})$ is directly sampled from a multivariate normal distribution, specifically $\operatorname{MVN}(\mathbf{0},\mathbf{I}_{4\times 4})$, where $\mathbf{I}_{4\times 4}$ denotes the diagonal unit matrix. Moreover, we generate the true parameters $\gamma^{*}$ and $\beta^{*}$ from a multivariate normal distribution $\operatorname{MVN}(\mathbf{0},\mathbf{I}_{4\times 4})$. We normalize the feature and true parameters to satisfy $\|\phi(a,X_{t})\|\leq 1$, $\|\gamma^{*}\|\leq 1$ and $\|\beta^{*}\|\leq 1$.
Our simulation studies encompass a total of eight settings, with the stochastic proximal outcome generated from (1) a Poisson distribution, (2) an overdispersed Poisson (OP) distribution with a low dispersion level, (3) an OP distribution with a moderate dispersion level, (4) an OP distribution with a high dispersion level, (5) a ZIP distribution, (6) a zero-inflated and overdispersed Poisson (ZIOP) distribution with a low dispersion level, (7) a ZIOP distribution with a moderate dispersion level, and (8) a ZIOP distribution with a high dispersion level.
@@ 6.1 Setup
$Y_{t}\sim\operatorname{Poisson}(\lambda_{t}\mu_{t}),~{}\ \lambda_{t}\sim\operatorname{Gamma}(\omega,1/\omega),~{}\ \mu_{t}=\exp\left(\phi(A_{t},X_{t})^{\top}\beta^{*}\right).$ (5)
In this case, the outcomes are generated from an OP model, where $\lambda_{t}$ introduces extra overdispersion, and $\omega$ controls the level of dispersion. We induce high, moderate, and low levels of overdispersion by setting $\omega=0.25,1,$ and $25$, respectively. When setting $\lambda_{t}=1$, we recover the Poisson distribution.
@@
$Y_{t}$ $=$ $Z_{t}C_{t}.$ (6)
Here, the outcomes are generated from a ZIOP model. We also consider scenarios without zero-inflation and overdispersion by setting both $Z_{t}$ and $\lambda_{t}$ to $1$.
In all simulations, we set the tuning parameter $\alpha=1$, which implements the Laplace approximation. We designated the initial random exploration period as $\tau=20$, during which each intervention was selected once to establish the initial estimators. Performance was assessed by averaging frequentist regret over 200 replications.
To further evaluate performance, we compare the proposed algorithms with their TS counterparts, which use MCMC methods to approximate posterior distributions. For MCMC-based TS algorithms, we use priors $\beta^{*}\sim\mathcal{N}(0,I)$ and $\gamma^{*}\sim\mathcal{N}(0,I)$. We assess the total running time and regret over $200$ replications. The stochastic outcomes are generated under Setting (7). Due to computational constraints, we set $T=200$ for these evaluations.
@@ 6.2 Results
The results under Settings (1) - (4) are presented in Figure 2. Our findings reveal that the TS-Count variants exhibit superior performance compared to the Linear TS algorithm even with log transformation, which demonstrates consistently high and linear regret across all settings. Furthermore, our results suggest that TS-Poisson outperforms other algorithms, including TS-NB, particularly in settings with varying degrees of overdispersion. This is attributed to the reliability of Poisson regression in yielding consistent estimates when the conditional mean model is appropriately specified, meeting the requirement for sublinear regret, as demonstrated in the proof of regret.
@@
In Figure 4, we present the comparison results between the proposed algorithms and their MCMC counterparts. Figure 4a demonstrates that the proposed TS-Count algorithms significantly reduce computation time compared to their MCMC counterparts. Figure 4b reveals that the proposed algorithms achieve similar levels of regret as their MCMC versions. Overall, the proposed algorithms offer substantial computational efficiency without compromising performance, making them practically advantageous.
@@
(a) Computation time
(b) Regret
Figure 4: Comparison of the proposed algorithms and their MCMC counterparts focuses on computation time and regret. The results are calculated from $200$ replications of the experiment.
@@ 7 Application to the Drink Less Case Study
In this section, we implement the proposed algorithms in a real case study - the Drink Less MRT - to evaluate and benchmark their performances. We consider not only the TS-Count variants and the Linear TS algorithm that incorporates a logarithmic transformation of the outcome but also the original static strategy from Drink Less, which assigns interventions at a constant probability $0.6$. Our application is divided into two parts: (1) a simulation study guided by the Drink Less data, and (2) an off-policy evaluation directly applied to the Drink Less data.
7.1 Implementation of Algorithms
@@
Up to this point, our discussion has centered on an individual user. However, within the framework of mHealth studies or MRTs, the paradigm typically involves $N$ users, each receiving interventions at $T$ decision times. We introduce the subscript $n$ to denote each user. At any decision time $t$, we have a collection of historical observations denoted by $D_{n,t}=\{X_{n,i},A_{n,i},Y_{n,i};i=1,\cdots,t-1\}_{n=1}^{N}\cup\{X_{n,t}\}_{n=1}^{N}$. The overall objective is to customize interventions to cater to the specific demands of each user. This scenario can be envisioned as $N$ separate contextual bandit problems, where each participant is associated with a unique model parameter $\eta_{n}^{*}$. We implement the proposed algorithms individually for each user, consistent with previous research that has shown the effectiveness of personalized approaches (Liao et al., 2020; Yao et al., 2021).
Recall that $\phi(A_{n,t},X_{n,t})$ represents the feature vector, which transforms the intervention $A_{n,t}$ and the context $X_{n,t}$ as follows:
$\phi(A_{n,t},X_{n,t})^{\top}=(X_{n,t}^{\top},A_{n,t}S_{n,t}^{\top}),$
where $X_{n,t}$ is the context containing baseline and time-varying information of user $n$, and $S_{n,t}$ is a subset of $X_{n,t}$ that may interact with intervention (also known as moderators in the causal inference literature). Table S1 in Section S4.1 of the Supplement describes the context available in the Drink Less study.
@@
Another critical aspect of applying the proposed algorithms to MRTs is ensuring robust post-study statistical analyses. It is important to note that algorithms designed solely to minimize regret may not ensure error control or provide sufficient statistical power to draw reliable inferences about treatment effects using standard methods, nor may they effectively support off-policy evaluations (Liao et al., 2020; Yao et al., 2021). To address this issue, the randomization probabilities must remain strictly between 00 and $1$, avoiding extremes. Specifically, we constrain the randomization probability within bounds $p_{\min}$ and $p_{\max}$, i.e., $\tilde{p}_{t}(X_{n,t})=\max(p_{\min},\min(p_{\max},p_{t}(X_{n,t}))$. This method is known as “clipping” in mHealth (Liao et al., 2020; Yao et al., 2021). For our simulations, we set $p_{\min}=0.01$ and $p_{\max}=0.99$.
Moreover, we apply the proposed algorithms to the Drink Less Data by incorporating a ridge penalty of the form $\lambda\|\beta_{n}-\beta_{0}\|_{2}^{2}$ and $\lambda\|\gamma_{n}-\gamma_{0}\|_{2}^{2}$, rather than incorporating an explicit initial exploration stage of length $\tau$. This modification is enabled by the availability of parallel data from the Drink Less MRT (Bell et al., 2020), allowing us to form prior estimates for $\beta_{n}^{*}$ and $\gamma_{n}^{*}$. Following the guidelines from Theorems 5.3 and 5.4, we choose $\lambda=1$. For subsequent analyses, $\beta_{0}$ and $\gamma_{0}$ are set based on the estimates from the parallel data.
7.2 Clipped Regret Analysis
Due to the implementation of clipping, the traditional regret metric is not applicable. Instead, we assess the proposed algorithms against a clipped oracle - an oracle constrained to intervention probabilities within the range $[p_{\min},p_{\max}]$ (Greenewald et al., 2017; Yao et al., 2021). Specifically, the frequentist regret relative to this clipped oracle is defined as follows:
$\mathcal{R}(T,\eta^{*})=\sum_{t=1}^{T}\mathbb{E}\left[\max_{a,p^{*}}h(\phi(a,X_{t})^{\top};\eta^{*})-h(\phi(A_{t},X_{t})^{\top};\eta^{*})\Big{|}\eta^{*}\right],$
where $a\in\mathcal{A}$ and $p^{*}\in[p_{\min},p_{\max}]$. It can be shown that, with probability clipping, the proposed TS-Count algorithms will maintain the same regret bounds with respect to a clipped oracle. Here, we consider TS-NB as a specific example. Suppose Let $\bar{A}_{t}^{*}$ represent the optimal non-zero intervention chosen by the oracle and $\bar{A}_{t}$ is the corresponding choice made by the algorithm. Following Greenewald et al. (2017), the regret for times $t>\tau$ can be decomposed as:
$\operatorname{regret}_{t}$ $=$ $p^{*}\exp(\phi(\bar{A}_{t}^{*},X_{t})^{\top}\beta^{*})+(1-p^{*})\exp(\phi(0,X_{t})^{\top}\beta^{*})$
$-\tilde{p}_{t}\exp(\phi(\bar{A}_{t},X_{t})^{\top}\beta^{*})-(1-\tilde{p}_{t})\exp(\phi(0,X_{t})^{\top}\beta^{*})$
$\leq$ $(p^{*}-\tilde{p}_{t})[\exp(\phi(\bar{A}_{t},X_{t})^{\top}\beta^{*})-\exp(\phi(0,X_{t})^{\top}\beta^{*})]$
$+$ $[\exp(\phi(\bar{A}_{t}^{*},X_{t})^{\top}\beta^{*})-\exp(\phi(\bar{A}_{t},X_{t})^{\top}\beta^{*})]$
Under Assumptions 1 - 3, the cumulative regret $\mathcal{R}(T,\beta^{*})$ can be bounded as follows:
$\mathcal{R}(T,\beta^{*})\leq\tau\Delta_{\max}+\sum_{t=\tau+1}^{T}\mathbb{E}[\operatorname{regret}_{t}]\leq\tau\Delta_{\max}+\sum_{t=\tau+1}^{T}\mathbb{E}[T_{1}+T_{2}]$
where $T_{1}=(p^{*}-\tilde{p}_{t})[\exp(\phi(\bar{A}_{t},X_{t})^{\top}\beta^{*})-\exp(\phi(0,X_{t})^{\top}\beta^{*})]$ and $T_{2}=[\exp(\phi(\bar{A}_{t}^{*},X_{t})^{\top}\beta^{*})-\exp(\phi(\bar{A}_{t},X_{t})^{\top}\beta^{*})]$. We bound $\mathbb{E}[T_{1}]$ based on Lemma 1 of Greenewald et al. (2017) and bound $\mathbb{E}[T_{2}]$ using Lemma S2.1 from the Supplement. According to Greenewald et al. (2017), $\mathbb{E}[T_{2}]$ is the predominant component affecting the overall regret bound. Consequently, the clipped frequentist regret bound for TS-NB aligns with the standard unclipped scenario. This reasoning is similarly applicable to other variants of the TS-Count algorithms.
7.3 Simulation Study Guided by Drink Less
In this part, we build a simulation environment to mimic the Drink Less MRT and evaluate the candidate algorithms. We show that the proposed algorithms incur lesser user-averaged regret.
7.3.1 Simulation Environment
We consider two different outcome generative models for the Drink Less study: (1) an OP model, as presented in Equation (5), and (2) a ZIOP model, as presented in Equation (6.1). We use the Drink Less MRT data to fit the outcome model. Each model was fitted individually for each user using the maximum a posteriori (MAP) estimation method. We used a prior $\beta_{n}^{*}\sim\mathcal{N}(0,I)$ for model (1) and an additional prior $\gamma_{n}^{*}\sim\mathcal{N}(0,I)$ for model (2). In our simulations, the context $X_{n,t}$ is directly observed from the data to mimic the practical scenario accurately.